Determinate
3y-z-4=0
Y+2z-13=0
2y-5z+11=0

Answer by stanbon(75887)

Determinate
3y-z-4=0
Y+2z-13=0
2y-5z+11=0
-------
z = 3y-4
---
Substitute into the 2nd equation and solve for "y":
y + 2*(3y-4)-13 = 0
7y - 21 = 0
y = 3
----
Therefore z = 3*3-4 = 5
----
Checking:
2y-5z+11 = 0
6-25+11 = 0
8 = 0
False
Conclusion: The system of 3 equations is inconsistent
===================
Cheers,
Stan H.
====================